OpenStudy (anonymous):

find q(t)=ab^t, if q(5)=5 and q(7)=45

OpenStudy (anonymous):

first plug t = 5 and get 5 = a * b^5 the plug t = 7 and get 45 = a * b^7 can you do it by yourself now ?

OpenStudy (anonymous):

here is a hint, what is \[\frac{ab^7}{ab^5}\] ?

OpenStudy (anonymous):

i did plug in those numbers, but i still cant know how to go further, i still need some assistance

OpenStudy (anonymous):

7/5?

OpenStudy (anonymous):

45 = a * b^7 5 = a * b^5 divide the first be the second what do you get ?

OpenStudy (anonymous):

\[\frac{ab^7}{ab^5}=b^2\] is one answer

OpenStudy (anonymous):

9= b^2

OpenStudy (anonymous):

very good now take the square root of it

OpenStudy (anonymous):

the other answer is \(\frac{45}{5}=9\) which tells you that \(b^2=9\) so you can find \(b\) once you have \(b\) substitute back to find \(a\)

OpenStudy (anonymous):

thanks guys, really appreciate the assistance